# Generalizable data-driven turbulence closure modeling on unstructured grids with differentiable physics

### v2

@@ 2.2 Graph neural networks
From the ansatz presented in the prior section, both the eddy viscosity and nonlinear contributions of the SGS stress can be estimated using a deep learning model. $\bm{\tau}^{EV}$ is a function of a single scalar coefficient $C_{s}$, while $\bm{\tau}^{NL}$ represents a rank-2 tensor whose components can be arbitrarily learned. We set a few requirements to determine a choice of DNN architecture. We wish for $C_{s}$ and $\bm{\tau}^{NL}$ to be both spatiotemporally varying and functions of the instantaneous resolved velocity and pressure fields $\overline{\mathbf{u}},\overline{p}$. In addition, it is understood that the SGS stress is non-local [66], in that the stress at a point in space is influenced not only by the local flow, but also the surrounding flow as well. Lastly, the network architecture must be amenable to the choice of discretization of the coarse-grained resolved solution field, which, to be most generalizable, can be assumed to be an unstructured computational mesh. Based on these requirements, we propose to model the desired turbulence quantities $C_{s}$ and $\tau^{NL}_{ij}$ using graph neural networks. Here, we use the Encode-Process-Decode paradigm that is commonly found in deep GNN models. We provide an overview of the GNN architecture in this section.
2.2.1 Graph representation of flowfield
@@
The input graph is defined by $G_{0}=(\mathbf{V_{0}},\mathbf{E_{0}},\mathbf{A})$, where $\mathbf{V_{0}},\mathbf{E_{0}}$ and $\mathbf{A}$ represent the node features, edge features and adjacent matrix between nodes, respectively. Specifically, $\mathbf{V_{0}}$ contains the feature vectors for each node, while $\mathbf{E_{0}}$ contains the feature vectors for each edge. The adjacency matrix $\mathbf{A}$ is a binary (one-hot) matrix that defines the connections between nodes, where $\mathbf{A}_{ij}=1$ if there is an edge between node $i$ and node $j$, and $\mathbf{A}_{ij}=0$ otherwise.
To represent flowfield data from an FEM discretization as a graph, all nodes in the finite element mesh are mapped to graph nodes, $\mathbf{V_{0}}$, while the edges of finite elements between corresponding nodes are represented by graph edges, $\mathbf{E_{0}}$. Additionally, extra edges are generated based on a radial cutoff, where connections between nodes are defined by a cutoff radius, the Euclidean distance, and the maximum number of neighbors. This procedure is briefly illustrated in Fig. 2. Nodes within the cutoff radius are sequentially added, starting from the one with the smallest Euclidean distance, until the total number of connected nodes reaches the specified maximum number of neighbors.
Figure 2: Schematic of generation of additional edges based on radial cutoff and definition of node and edge features; maximum number of neighbors = 8
@@
The node features, $\mathbf{V_{0}}$, are composed of both fixed and time-varying components. The fixed node features include the position vector, the binary wall indicator, and the characteristic length scale associated with each node. The position vector of each node, $\mathbf{x}_{i}$, defines the spatial arrangement of the computational mesh, providing a point cloud on which the required SGS stress model parameters are computed. This feature is time-invariant because the mesh coordinates remain fixed. The binary wall indicator, $\delta_{\text{wall}}$, identifies whether a node is located on a no-slip wall ($\delta_{\text{wall}}=1$) or not ($\delta_{\text{wall}}=0$). This feature is crucial for near-wall modeling, as SGS stress is significantly influenced by boundary conditions at walls. The characteristic length scale, $\Delta_{\text{node}}$, is the same quantity used to compute the eddy viscosity contribution in the SGS stress model. It is calculated as the minimum edge length of each element at the cell center, and then interpolated to the nodes to define $\Delta_{\text{node}}$. The time-varying node features consist of the local resolved velocity gradient tensor, $\nabla\overline{\mathbf{u}}_{i}$, and the resolved pressure field, $\overline{p}_{i}$. A basis transformation is applied to the velocity gradient tensor to decompose it into its isotropic, symmetric-traceless, and skew-symmetric components. The edge features, $\mathbf{E_{0}}$, include the relative position vector between nodes, $\mathbf{r}_{ij}=\mathbf{x}_{i}-\mathbf{x}_{j}$, which is time-invariant since the node positions are fixed. The node and edge features are briefly described in Fig. 2. We remark that time-varying FEM discretizations are also compatible with graph representations but are not implemented in this study.
2.2.2 Encode-Process-Decode
@@
We adopt a simple Encode-Process-Decode architecture for our GNN implementation as follows. After the graph is generated from the raw dataset, the node features in the input graph, $\mathbf{V_{0}}$, are mapped to a higher-dimensional latent node space, $\mathbf{h}_{i}$, through a linear projection. Similarly, the latent edge features, $\mathbf{h}_{ij}$, are obtained by applying a linear projection to the edge features from the input graph. The encoding process is formally defined as follows:
@@
$\mathbf{h}_{ij}$ $=\mathcal{E}_{e}(\mathbf{E_{0}})$ (9)
where $\mathcal{E}_{v}$ and $\mathcal{E}_{e}$ are linear projection-based encoders for node and edge features with the structure of $[9,128]$ and $[2,128]$, respectively, and are applied node/edge-wise. Once the input graph’s node and edge features are encoded into latent node and edge representations, the processor performs multiple rounds of message passing on these latent features. In our setup, we use three message-passing layers within the processor. Each layer consists of arithmetic operations and multi-layer perceptrons (MLPs), with batch normalization applied after each message-passing step to stabilize learning.
After processing, the node features are decoded using a linear projection to predict the target quantities for the SGS stress model. Our model predicts a total of four outputs in two dimensions (2D): the SGS coefficient $C_{s}$ and four independent components of the non-linear stress tensor $\tau^{NL}_{ij}$, where $i$ and $j$ refers to the Cartesian components of a symmetric 2D tensor. Specifically, the isotropic and symmetric-traceless components of $\tau^{NL}_{ij}$ are predicted separately and then summed to form the final stress tensor, which is defined as follows:
$\bm{\tau}^{NL}$ $=k\begin{bmatrix}1&0\\ 0&1\end{bmatrix}+\alpha\begin{bmatrix}1&0\\ 0&-1\end{bmatrix}+\beta\begin{bmatrix}0&1\\ 1&0\end{bmatrix}+\gamma\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$ (10)
Then, the decoding process is defined as follows:
$\alpha_{\bm{\tau}}$ $=\mathcal{D}_{v}(\mathbf{h}_{i}),$ (11)
where $\alpha_{\bm{\tau}}$ is predicted turbulence model parameters including $(C_{s},k,\alpha,\beta,\gamma)$, and $\mathcal{D}_{v}$ is a linear projection-based decoder for the processed node features $\mathbf{h}_{i}$, with structure of $[128,5]$.
2.2.3 Message-passing
@@
where $f_{e}$ is an edge update function containing MLPs with the structure of $[384,128,128]$. The sigmoid linear unit (SiLU) serves as the activation function at the hidden layer.
@@
$\mathbf{h}_{i}$ $=\mathbf{h}_{i}+\mathbf{h}_{i}^{\prime}$ (15)
where $\mathcal{N}$ represents a neighborhood of connectivity and $f_{v}$ is a node update function containing trainable MLPs with the structure of $[256,128,128]$ and SiLU serving as the activation function at the hidden layer.
2.3 Implementation of the adjoint method
The object of the differentiable turbulence is to pair the deep learning model with a differentiable solution algorithm for the governing equations, such that the parameters of the network can be trained in an end-to-end fashion. Automatic differentiation (AD) frameworks have become a necessary tool for training deep neural networks using gradient-based optimization. The AD principle relies on the fact that gradients of a composition of differentiable functions can be exactly computed using the chain rule. Most standard DNNs are composed of straightforward feed-forward operations, such as linear operations and pointwise nonlinearities, which can easily be written in high-level scientific programming languages. However, the integration of PDE solution algorithms, particularly using unstructured computational meshes, can add enormous complexity that makes writing these algorithms in an AD-enabled framework impractical. While the development of PDE solvers in AD-enabled languages is underway, there currently exists a plethora of efficient and robust non-differentiable PDE solvers that can interface with high-level languages and be called as external routines. If these can be integrated within an AD framework such as PyTorch [67], we can leverage the advantages of a differentiable solution scheme without rewriting complex PDE solvers from scratch.
@@ 2.3 Implementation of the adjoint method
The iterative nature of Eq. 19 allows gradients of $y$ to be propagated backwards starting from $\frac{dy}{dx_{n}}$. For each step in the forward pass, Eqs. 16-18, a corresponding reverse function should be implemented, also known as a vector-jacobian product (VJP):
@@
Initialize:
      GNNθ
tentative_vel
pres_correct
vel_correct
$\mathbf{u},p$
$t\leftarrow 0$
while $t<T$ do
     $\alpha_{\bm{\tau}}\leftarrow\texttt{GNN}_{\theta}(\mathbf{u},p)$
     $\mathbf{u}^{*}\leftarrow\texttt{tentative\_vel}(\mathbf{u},p,\alpha_{\bm{\tau}})$
     $p^{*}\leftarrow\texttt{pres\_correct}(\mathbf{u}^{*},p)$
     $\mathbf{u}\leftarrow\texttt{vel\_correct}(\mathbf{u}^{*},p^{*},p)$
     $p\leftarrow p^{*}$
     $t\leftarrow t+\Delta t$
end while
The three PDE-solving functions can be represented in the FeniCS formalism as:
@@
$y_{\mathbf{m}}=-\bm{\lambda}^{T}\left(\mathbf{A}_{\mathbf{m}}-\mathbf{b}_{\mathbf{m}}\right),$ (24)
@@
Our choice of FEniCS [58] for this study is motivated by this criteria. FEniCS offers a general-purpose PDE solving framework using the finite element method. FEniCS provides the ability to define PDEs symbolically using unified form language (UFL) [68] and automatic differentiation functionality that is useful for computing $\mathbf{A}_{\mathbf{m}}$ and $\mathbf{b}_{\mathbf{m}}$. In addition, FEniCS incorporates an integrated PETSc library [69], which facilitates the efficient computation of linear systems during the solution of PDEs and the discrete adjoint method. PDEs are defined using their variational form. For example, a Poisson system can be defined as: {python} F = inner(nu*grad(u), grad(v))*dx where u is a TrialFunction, v is a TestFunction, and nu is a given Function. The matrix $\mathbf{A}$ at the left-hand side and vector $\mathbf{b}$ at the right-hand side of the discretized linear system can be produced by calling the assemble routine on F. To compute $(\mathbf{A}-\mathbf{b})_{\mathrm{\texttt{nu}}}$, one need only assemble a new form: {python} F_nu = derivative(F, nu) With these procedures in hand, a custom VJP for PDE_solve can be supplied, described in Algorithm 2.
@@
function PDE_solve($\mathbf{m};$ F)
     $\mathbf{A},\mathbf{b}=$ assemble(F(m)) $\triangleright$ LHS/RHS of form F
     $\mathbf{x}=$ solve($\mathbf{A},\mathbf{b}$) $\triangleright$ Linear solve
     return x
end function
function PDE_solve_vjp($y_{\mathbf{x}},\mathbf{m};$ F)
     $\bm{\lambda}=$ solve($-\mathbf{A}^{T},y_{\mathbf{x}}^{T}$) $\triangleright$ Adjoint solution
     $F_{\mathbf{m}}=$ derivative(F(m; x), m) $\triangleright$ Residual gradient
     $y_{\mathbf{m}}=$ assemble($-\bm{\lambda}^{T}F_{\mathbf{m}}$) $\triangleright$ Assembled gradient
     return $y_{\mathbf{m}}$
end function
Note that for brevity, we have included only the major steps of the procedure and removed many of the specific details of the implementation, such as conversion of PyTorch tensors to FEniCS variables, the application of boundary conditions, and certain FEniCS syntax. With PDE_solve_vjp implemented, PDE_solve can be used as any other differentiable PyTorch layer, and the entire IPCS algorithm can be written using PyTorch variables, which naturally record the gradient tape required for automatic backwards differentiation.
3 Datasets and training
@@
We first describe our training configuration given by flow over a backwards-facing step (BFS). This is a common benchmark problem in fluid dynamics that has been widely studied in both experimental [73] and computational [74] settings. BFS flow is characterized by flow separation at the step and a subsequent reattachment region further downstream, which introduces a recirculation region near the step and transient vortex shedding. These phenomena are found frequently in more complex non-idealized flows and can be challenging to model accurately, making the BFS a valuable case study for evaluating modeling approaches. The BFS configuration can be described as flow in a channel with a sudden expansion. An inlet is prescribed on the left with a uniform velocity in the x-direction, entering a channel of fixed width. Once the flow has fully developed, the channel width is increased, creating a step at the lower wall. This expansion produces a shear layer at the step height, which leads to turbulent mixing and vortex development. The flow reattaches downstream and eventually exits the domain through a pressure outlet. A schematic of the computational domain is presented in Fig. 3(a).
@@
Figure 3: (a) Schematic of the backwards-facing step domain with inlet, outlet, and dimensions labeled. (b) A subsection of the “5k” mesh used to evaluate the ML closure. Greater density of elements is placed in front of the step.
@@
Ground truth data is taken from a high-resolution LES simulation of the BFS flow under the Reynolds number, $Re=U_{\infty}h_{0}/\nu=26,051$, where $U_{\infty}=32m/s$ and $\nu=1.56\times 10^{-5}$. We use an unstructured triangular mesh of on average 80,000 Taylor-Hood elements, which is refined near all walls and particularly around and downstream from the step. The highest density of elements resides in the shear layer at the step, with generally decreasing density further away. The LES equations using the standard Smagorinsky turbulence model with $C_{s}=0.16$ are solved using an implicit pressure correction scheme and Crank-Nicolson timestepping. We additionally use streamline upwind Petrov-Galerkin (SUPG) and least square incompressibility constraint (LSIC) stabilization due to the finite element implementation. For consistency, we set the ground truth timestep $\Delta t$ to $5\times 10^{-6}s$, which is sufficient to keep the Courant-Friedrichs-Lewy (CFL) number approximately equal to or less than 1, where $\mathrm{CFL}<1$ is not strictly required due to the implicit timestepping. Snapshots of the velocity and pressure fields are saved every $10^{-4}$ seconds, and a total of 1024 snapshots are generated. We denote results from this ground truth simulation as “LES-80k”. More comprehensive efforts to align the modeling with experimental or direct numerical simulation (DNS) studies under specific operating conditions are considered beyond the scope of this work, because the primary objective of this work is to demonstrate the capabilities of the graph neural network and the differentiable framework.
@@
We generate a significantly coarser mesh for the BFS flow, consisting of 5,253 elements, on which we apply our ML-based closure and evaluate it against baseline comparisons. Fig. 3(b) shows a portion of this mesh in the area around the step. The scaling factor for element sizes was set to be 4.5$\approx$ 4.5$\times$ 4.5 × the size of the high-resolution mesh. Simulations on this coarse grid use nearly the same solution scheme as the ground truth, except the SUPG stabilization is removed and the timestep size is increased to $\Delta t=10^{-4}s$. We test two baseline comparisons on this grid, no model (“NM-5k”), where no closure model is used, and the standard Smagorinsky closure (“LES-5k”), where we set $C_{s}=0.06$ based on optimizing a-posteriori results with respect to the ground truth. The ML-closure is denoted “ML-5k”.
@@
A third mesh, consisting of 23,502 elements, is also created for each geometry, positioned between the coarse ML mesh and the fine ground truth mesh in terms of size. This mesh serves as an additional baseline comparison (referred to as ’LES-20k’) for the model in relation to a more well-resolved LES simulation. The Smagorinsky model is used with $C_{s}=0.07$, and the timestep is reduced to $\Delta t=5\times 10^{-5}s$.

### v3

@@ II.2 Graph neural networks
From the ansatz presented in the prior section, the SGS stress can be estimated using a deep learning model. We set a few requirements to determine a choice of DNN architecture. We wish for $C_{s}$ to be spatiotemporally varying and a function of the instantaneous resolved velocity and pressure fields $\overline{\mathbf{u}},\overline{p}$. In addition, it is understood that the SGS stress is non-local [75], in that the stress at a point in space is influenced not only by the local flow, but also the surrounding flow as well. Lastly, the network architecture must be amenable to the choice of discretization of the coarse-grained resolved solution field, which, to be most generalizable, can be assumed to be an unstructured computational mesh. Based on these requirements, we propose to model the desired turbulence quantity, $C_{s}$, using graph neural networks (GNNs). Here, we use the Encode-Process-Decode paradigm that is commonly found in deep GNN models. We provide an overview of the GNN architecture in this section.
II.2.1 Graph representation of flowfield
@@
The input graph is defined by $G_{0}=(\mathbf{V_{0}},\mathbf{E_{0}},\mathbf{A})$, where $\mathbf{V_{0}},\mathbf{E_{0}}$ and $\mathbf{A}$ represent the node features, edge features and redadjacency matrix between nodes, respectively. Specifically, $\mathbf{V_{0}}$ contains the feature vectors for each node, while $\mathbf{E_{0}}$ contains the feature vectors for each edge. The adjacency matrix $\mathbf{A}$ is a binary (one-hot) matrix that defines the connections between nodes, where $\mathbf{A}_{ij}=1$ if there is an edge between node $i$ and node $j$, and $\mathbf{A}_{ij}=0$ otherwise.
@@
To represent flowfield data from an FEM discretization as a graph, all nodes in the finite element mesh are mapped to graph nodes, $\mathbf{V_{0}}$, while the edges of finite elements between corresponding nodes are represented by graph edges, $\mathbf{E_{0}}$, as in Fig. 2. The node features, $\mathbf{V_{0}}$, are composed of both fixed and time-varying components. The fixed node features include the position vector, the binary wall indicator, and the characteristic length scale associated with each node. The position vector of each node, $\mathbf{x}_{i}$, defines the spatial arrangement of the computational mesh, providing a point cloud on which the required SGS stress model parameters are computed. This feature is time-invariant because the mesh coordinates remain fixed. The binary wall indicator, $\delta_{\text{wall}}$, identifies whether a node is located on a no-slip wall ($\delta_{\text{wall}}=1$) or not ($\delta_{\text{wall}}=0$). This feature is crucial for near-wall modeling, as SGS stress is significantly influenced by boundary conditions at walls. The characteristic length scale, $\Delta_{\text{node}}$, is the same quantity used to compute the eddy viscosity contribution in the SGS stress model. It is calculated as the minimum edge length of each element at the cell center, and then interpolated to the nodes to define $\Delta_{\text{node}}$. The time-varying node features consist of the local resolved velocity gradient tensor, $\nabla\overline{\mathbf{u}}_{i}$, and the resolved pressure gradient vector, $\nabla\overline{p}_{i}$. A basis transformation is applied to the velocity gradient tensor to decompose it into its isotropic, symmetric-traceless, and skew-symmetric components. The edge features, $\mathbf{E_{0}}$, include the relative position vector between nodes, $\mathbf{r}_{ij}=\mathbf{x}_{i}-\mathbf{x}_{j}$, which is time-invariant since the node positions are fixed. As a result, for the two-dimensional BFS case, the node and edge feature dimensions are 10 and 2; for the three-dimensional case, they are 17 and 3, respectively. The node and edge features for the two-dimensional BFS case are briefly described in Fig. 2. We remark that time-varying FEM discretizations are also compatible with graph representations but are not implemented in this study.
Figure 2: Schematic of GNN nodes and edges representing the FEM-based computational mesh for the two-dimensional BFS case.
II.2.2 Encode-Process-Decode
@@
We adopt a simple Encode-Process-Decode architecture for our GNN implementation as follows. After the graph is generated from the raw dataset, the node features in the input graph, $\mathbf{V_{0}}$, are mapped to a higher-dimensional latent node space, $\mathbf{h}_{i}$, through a linear projection. Similarly, the latent edge features, $\mathbf{h}_{ij}$, are obtained by applying a linear projection to the edge features from the input graph. Here, we set the dimension of latent node and edge spaces to 128 for the two-dimensional case. For the three-dimensional case, we increase the latent dimension to 256 to accommodate the larger node/edge feature dimensionality (17/3 vs. 10/2) and the greater complexity of three-dimensional turbulence, which requires higher representational capacity. This enrichment of node and edge feature representations enhances the expressiveness of the architecture, enabling subsequent GNN layers to more effectively capture nonlinear relationships within local graph neighborhoods. The encoding process is formally defined as follows:
@@
$\mathbf{h}_{ij}$ $=\mathcal{E}_{e}(\mathbf{E_{0}}),$ (8)
where $\mathcal{E}_{v}$ and $\mathcal{E}_{e}$ are linear projection-based encoders for node and edge features with the structure of $[17,256]$ and $[3,256]$ for the three-dimensional case ($[10,128]$ and $[2,128]$ for 2D), respectively, and are applied node/edge-wise. This reflects the input feature sizes (see Sec. II.2.1.) and the latent widths used (256 in 3D; 128 in 2D). Once the input graph’s node and edge features are encoded into latent node and edge representations, the processor performs multiple rounds of message passing on these latent features. In our setup, we use four message-passing layers within the processor. Each layer consists of arithmetic operations and multi-layer perceptrons (MLPs).
After processing, the node features are decoded using a linear projection to predict the target quantity, $C_{s}$, for the SGS stress model. Then, the decoding process is defined as follows:
$C_{s}$ $=\mathcal{D}_{v}(\mathbf{h}_{i}),$ (9)
where $\mathcal{D}_{v}$ is a linear projection-based decoder for the processed node features $\mathbf{h}_{i}$, with the structure of $[256,1]$ for the three-dimensional case ($[128,1]$ for 2D).
II.2.3 Message-passing
@@
where $f_{e}$ is an edge update function containing MLPs with the structure of $[768,256,256]$ for the three-dimensional case ($[384,128,128]$ for 2D). The sigmoid linear unit (SiLU) serves as the activation function at the hidden layer.
@@
$\mathbf{h}_{i}$ $=\mathbf{h}_{i}+\mathbf{h}_{i}^{\prime},$ (13)
where $\mathcal{N}$ represents a neighborhood of connectivity and $f_{v}$ is a node update function containing trainable MLPs with the structure of $[768,256,256]$ for the three-dimensional case ($[256,128,128]$ for 2D) and SiLU serving as the activation function at the hidden layer.
II.3 Implementation of the adjoint method
The objective of differentiable physics based turbulence closure modeling is to pair the deep learning model with a differentiable solution algorithm for the governing equations, such that the parameters of the neural network predicting a closure model can be trained in an end-to-end fashion. Automatic differentiation (AD) frameworks have become a necessary tool for training deep neural networks using gradient-based optimization. In AD, compositions of differentiable functions can be exactly computed using the chain rule. Most standard DNNs are composed of straightforward feed-forward operations, such as linear operations and pointwise nonlinearities, which can easily be written in high-level scientific programming languages. However, the integration of PDE solution algorithms, particularly using unstructured computational meshes, can add enormous complexity that makes writing these algorithms in an AD-enabled framework impractical. While the development of PDE solvers in AD-enabled languages is underway [76], there currently exists a plethora of efficient and robust non-differentiable PDE solvers equipped with discrete adjoint solvers, that can interface with high-level languages and be called as external routines. If these can be integrated within an AD framework such as PyTorch [77], we can leverage the advantages of a differentiable solution scheme without rewriting complex PDE solvers from scratch.
@@ II.3 Implementation of the adjoint method
The iterative nature of Eq. 17 allows gradients of $y$ to be propagated backwards starting from $\frac{dy}{dx_{n}}$. For each step in the forward pass, Eqs. 14-16, a corresponding reverse function should be implemented, also known as a vector-Jacobian product (VJP):
@@
1
2Initialize:
3   GNNθ
4   pres_correct
5   vel_correct
6   $\mathbf{u},p$
7   $t\leftarrow 0$
8 for $t<T$ do
9  $\alpha_{\bm{\tau}}\leftarrow\texttt{GNN}_{\theta}(\mathbf{u},p)$
10  $\mathbf{u}^{*}\leftarrow\texttt{tentative\_vel}(\mathbf{u},p,\alpha_{\bm{\tau}})$
11  $p^{*}\leftarrow\texttt{pres\_correct}(\mathbf{u}^{*},p)$
12  $\mathbf{u}\leftarrow\texttt{vel\_correct}(\mathbf{u}^{*},p^{*},p)$
13  $p\leftarrow p^{*}$
14  $t\leftarrow t+\Delta t$
15
@@ 14  $t\leftarrow t+\Delta t$
The three PDE-solving functions can be represented in the FEniCS/FEniCSx formalism as:
@@
$y_{\mathbf{m}}=\bm{\lambda}^{T}\left(\mathbf{A}_{\mathbf{m}}\mathbf{x}-\mathbf{b}_{\mathbf{m}}\right),$ (22)
@@
Our choice of FEniCS/FEniCSx [70, 71] for this study is motivated by this criteria. FEniCS/FEniCSx offers a general-purpose PDE solving framework using the finite element method. FEniCS/FEniCSx provides the ability to define PDEs symbolically using unified form language (UFL) [78] and automatic differentiation functionality that is useful for computing $\mathbf{A}_{\mathbf{m}}$ and $\mathbf{b}_{\mathbf{m}}$. In addition, FEniCS/FEniCSx incorporates an integrated PETSc library [79], which facilitates the efficient computation of linear systems during the solution of PDEs and the discrete adjoint method. PDEs are defined using their variational form. For example, a Poisson system can be defined as:
where u is a TrialFunction, v is a TestFunction, and nu is a given Function. The matrix $\mathbf{A}$ at the left-hand side and vector $\mathbf{b}$ at the right-hand side of the discretized linear system can be produced by calling the assemble routine on F. To compute $(\mathbf{A}-\mathbf{b})_{\mathrm{\texttt{nu}}}$, one need only assemble a new form:
With these procedures in hand, a custom VJP for PDE_solve can be supplied, described in Algorithm 2.
1function PDE_solve($\mathbf{m};$ F)
2   $\mathbf{A},\mathbf{b}=$ assemble(F(m)) : LHS/RHS of form F
3   $\mathbf{x}=$ solve($\mathbf{A},\mathbf{b}$) : Linear solve
4   return x
5 end function
6
7function PDE_solve_vjp($y_{\mathbf{x}},\mathbf{m};$ F)
[…]
9   $F_{\mathbf{m}}=$ derivative(F(m; x), m) : Residual gradient
10   $y_{\mathbf{m}}=$ assemble($\bm{\lambda}^{T}F_{\mathbf{m}}$) : Assembled gradient
11   return $y_{\mathbf{m}}$
12 end function
@@ 10   $y_{\mathbf{m}}=$ assemble($\bm{\lambda}^{T}F_{\mathbf{m}}$) : Assembled gradient
Note that for brevity, we have included only the major steps of the procedure and removed many of the specific details of the implementation, such as conversion of PyTorch tensors to FEniCS/FEniCSx variables, the application of boundary conditions, and certain FEniCS/FEniCSx syntax. With PDE_solve_vjp implemented, PDE_solve can be used as any other differentiable PyTorch layer, and the entire IPCS algorithm can be written using PyTorch variables, which naturally record the gradient tape required for automatic backwards differentiation.
III Results and Discussion
III.1 Preliminary 2D validation
The purpose of this section is to demonstrate the feasibility of our framework using two-dimensional flow cases.
III.1.1 Dataset generation for 2D flows
@@
We first describe our training configuration given by flow over a backwards-facing step (BFS). This is a common benchmark problem in fluid dynamics that has been widely studied in both experimental [83] and computational [84] settings. BFS flow is characterized by flow separation at the step and a subsequent reattachment region further downstream, which introduces a recirculation region near the step and transient vortex shedding. These phenomena are found frequently in more complex non-idealized flows and can be challenging to model accurately, making the BFS a valuable case study for evaluating modeling approaches. The BFS configuration can be described as flow in a channel with a sudden expansion. An inlet is prescribed on the left with a uniform velocity in the $x$-direction, entering a channel of fixed width. Once the flow has fully developed, the channel width is increased, creating a step at the lower wall. This expansion produces a shear layer at the step height, which leads to turbulent mixing and vortex development. The flow reattaches downstream and eventually exits the domain through a pressure outlet. A schematic of the computational domain for the two-dimensional BFS is presented in Fig. 3(a).
@@
Figure 3: (a) Schematic of the two-dimensional backwards-facing step domain with inlet, outlet, and dimensions labeled. (b) A subsection of the “5k” mesh used to evaluate the ML closure. Greater density of elements is placed in front of the step.
@@
Ground truth data is taken from a high-resolution two-dimensional LES simulation of the BFS flow under the Reynolds number, $Re=U_{\infty}h_{0}/\nu=26,051$, where $U_{\infty}=32m/s$ and $\nu=1.56\times 10^{-5}$. We use an unstructured triangular mesh of on average 80,000 Taylor-Hood elements, which is refined near all walls and particularly around and downstream from the step. The elements with the highest density reside in the shear layer at the step, with generally decreasing density further away. The LES equations using the standard Smagorinsky turbulence model with $C_{s}=0.16$ are solved using an implicit pressure correction scheme and Crank-Nicolson timestepping. We employ only the least-squares incompressibility constraint (LSIC) stabilization in the finite element formulation, while the streamline-upwind Petrov-Galerkin (SUPG) term is deactivated. For consistency, we set the ground truth timestep $\Delta t$ to $5\times 10^{-6}s$, which is sufficient to keep the Courant-Friedrichs-Lewy (CFL) number approximately equal to or less than 1, where $\mathrm{CFL}<1$ is not strictly required due to the implicit timestepping. Snapshots of the velocity and pressure fields are saved every $10^{-4}$ seconds, and a total of 1024 snapshots are generated. We denote results from this ground truth simulation as “LES-80k”. More comprehensive efforts to align the modeling with experimental or DNS studies under specific operating conditions are considered beyond the scope of this work, because the primary objective of this work is to demonstrate the capabilities of the graph neural network and the differentiable framework.
@@
We generate a significantly coarser mesh for the two-dimensional BFS flow, consisting of 5,253 triangular cells with Taylor-Hood elements, on which we apply our ML-based closure and evaluate it against baseline comparisons. Fig. 3(b) shows a portion of this mesh in the area around the step. The scaling factor for element sizes was set to be approximately 4.5 times of the size of the high-resolution mesh. Simulations on this coarse grid use nearly the same solution scheme as the ground truth, except the timestep size is increased to $\Delta t=10^{-4}s$. We test two baseline comparisons on this grid, no model (“NM-5k”), where no closure model is used, and the standard Smagorinsky closure (“LES-5k”), where we set $C_{s}=0.06$ based on optimizing a-posteriori results with respect to the ground truth. The ML-closure is denoted “ML-5k”.
@@
A third mesh, consisting of 23,502 elements, is also created, positioned between the coarse ML mesh and the fine ground truth mesh in terms of size. This mesh serves as an additional baseline comparison (referred to as ’LES-20k’) for the model in relation to a more well-resolved LES simulation. The Smagorinsky model is used with $C_{s}=0.07$, and the timestep is reduced to $\Delta t=5\times 10^{-5}s$.
III.1.2 Training procedure and loss formulation